OpenStudy (anonymous):

h(x) = sin2 x + cos x 0 < x < 2π

OpenStudy (anonymous):

hello, is this for calculas?

OpenStudy (anonymous):

yes I need help with understanding derivatives for cofunctions.

OpenStudy (anonymous):

Calculus I

OpenStudy (anonymous):

ok r u looking for h'(x)?

OpenStudy (anonymous):

Yes. So the concept here is finding derivatives. In this situation product rule wouldn't apply correct?

OpenStudy (anonymous):

cos 2x + -sin x

OpenStudy (anonymous):

would this be correct?

OpenStudy (anonymous):

one sec

OpenStudy (kc_kennylau):

Nope, don't forget that you have to multiply the derivative of the substituted in the chain rule

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

f'(x) = cos(2x) (2) f'(x) = 2cos(2x)

OpenStudy (anonymous):

same for cosx

OpenStudy (kc_kennylau):

Wait, is it sin^2(x) or sin(2x)

OpenStudy (kc_kennylau):

The cos x part is correct though

OpenStudy (anonymous):

sin(2x)

OpenStudy (kc_kennylau):

Don't forget to multiply the derivative of 2x after applying the chain rule

OpenStudy (anonymous):

so f'(x)= 2cos(2x)-sinx

OpenStudy (kc_kennylau):

For example, [sin(5x)]' = 5cos(5x)

OpenStudy (anonymous):

oh gotcha. this is true

OpenStudy (kc_kennylau):

Yep, that's perfect

OpenStudy (anonymous):

thanks a lot. This helped out a lot. Ok so now how to convert this into radians.

OpenStudy (kc_kennylau):

You don't have to convert this into radians, because this was never degrees

OpenStudy (kc_kennylau):

remember, x is a variable

OpenStudy (anonymous):

yes true. I need to find the critical numbers though and that requires conversion. Correct?

OpenStudy (kc_kennylau):

What are you required to find?

OpenStudy (anonymous):

\[\frac{ d }{dx}\sin(2x)+\cos(x),(0<x<2\pi)\rightarrow 2\cos(2x)-\sin(x)!\]

OpenStudy (kc_kennylau):

Don't put the exclamation mark at the end, it is ambiguous to factorial

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

listen to @kc_kennylau

OpenStudy (anonymous):

there are two critical numbers for this equation between 0<theta<2pi

OpenStudy (kc_kennylau):

Set this to zero

OpenStudy (kc_kennylau):

That means solve h'(x)=0

OpenStudy (anonymous):

gotcha. Yes I remember now. Well thanks again.

OpenStudy (kc_kennylau):

no problem

OpenStudy (kc_kennylau):

You know how to solve the equation?

OpenStudy (anonymous):

This should be helpful

OpenStudy (anonymous):

I love khanacademy

OpenStudy (anonymous):

GTG, HAPPY STUDYING!!!!!